# FRASA: An End-to-End Reinforcement Learning Agent for Fall Recovery and Stand Up of Humanoid Robots

### v2

@@ IV-D Reward
[…]
@@ V-D Sim-to-real Transfer
Trunk CoM position offset (0, 0, 0) ±$\pm 5\times 10^{-3}$ m
@@
Actuator position ($q$) offset 0 ±$\pm 1.5$°
@@ V-F Fall Recovery Experiment
The length $L$ of the pendulum and the height $H$ of the stand are respectively $1.28$m and $1.74$m and the mass $m$ of the weight is $0.9$kg. Assuming the pendulum mass is point-like, the kinetic energy of the mass at the moment of impact varies from $4$J to $7.3$J across the different tested configurations, which are chosen to produce impacts ranging from mild disturbances to guaranteed falls. During the experiments, the velocity of the mass is zero after impact, indicating a complete transfer of energy to the robot.
@@ VI-B Quantitative results
In all configurations, FRASA achieves comparable or shorter instability times than the KFB method. The only configuration where the KFB method has a lower mean instability time than FRASA is for a $5.5$J disturbance from the back. However, the standard deviation ranges for the two methods overlap, indicating that the difference is not statistically significant.
For the most significant impacts ($d=0.89$m representing $7.3$J), FRASA is able to reject the disturbance in $42\%$ of the time required by the KFB method for a front impact and in $51\%$ of the time for a back impact. FRASA also demonstrates superior performance for $5.5$J frontwards disturbance, where the return to stability takes only $40\%$ of the time compared to the KFB method.

### v3

@@ IV-D Reward
No self-collision reward $R_{\text{collision}}=w_{3}\times\exp(-\text{self\_collisions})$
@@ V-D Sim-to-real Transfer
Trunk CoM position offset (0, 0, 0) $\pm 5\times 10^{-3}$ m
@@
Actuator position ($q$) offset 0 $\pm 1.5$°
@@ V-F Fall Recovery Experiment
The length $L$ of the pendulum and the height $H$ of the stand are respectively 1.281.28m and 1.741.74m and the mass $m$ of the weight is 0.90.9kg. Assuming the pendulum mass is point-like, the kinetic energy of the mass at the moment of impact varies from 44J to 7.37.3J across the different tested configurations, which are chosen to produce impacts ranging from mild disturbances to guaranteed falls. During the experiments, the velocity of the mass is zero after impact, indicating a complete transfer of energy to the robot.
@@ VI-B Quantitative results
In all configurations, FRASA achieves comparable or shorter instability times than the KFB method. The only configuration where the KFB method has a lower mean instability time than FRASA is for a 5.55.5J disturbance from the back. However, the standard deviation ranges for the two methods overlap, indicating that the difference is not statistically significant.
For the most significant impacts (d=0.89d=0.89m representing 7.37.3J), FRASA is able to reject the disturbance in $42\%$ of the time required by the KFB method for a front impact and in $51\%$ of the time for a back impact. FRASA also demonstrates superior performance for 5.55.5J frontwards disturbance, where the return to stability takes only $40\%$ of the time compared to the KFB method.